A corporation has 50,000 shares of $25 par stock outstanding that has a current market value of $150 per share. If the corporation issues a 5-for-1 stock split, the market value of the stock after the split will be approximately
Answer #1

$ 30

5 for 1 means for every shares, company will give five shares under stock split.

Stock split does not effect the market capitalization of shares. It effects only the quantity of shares. Now the outstanding shares after stock split = 50000*5

= 250000

Total market capitalization earlier = 50000*150

= 7500000

market value per share after split up = 7500000/250000 (market cap /shares after split up)

= 30

Please check with your answer and let me know.